OpenStudy (vortish):

find the minimum y-value on the graph of y=f(x) f(x)=-x^2+4x+8

OpenStudy (anonymous):

Have you taken calculus yet or are you in pre-calc/algebra?

OpenStudy (vortish):

algebra yes no cal

OpenStudy (anonymous):

Okay. Do you know how to complete the square? Doing that, you can rewrite the parabola in vertex form.

OpenStudy (vortish):

i understand some of it but i usually screw some thing up in it this is very hard for me i have never been any good at higher math

OpenStudy (anonymous):

\[\begin{align*}f(x)&=-x^2+4x+8\\ &=-\left(x^2-4x-8\right)\\ &=-\left(x^2-4x+4-12\right)\\ &=-\left((x-2)^2-12\right)\\ &=(x-2)^2+12\end{align*}\] So now f(x) is in vertex form, \[a(x-h)^2+k, \text{ where $(h,k)$ is the vertex.}\]

OpenStudy (vortish):

ok i understand that part of it

OpenStudy (anonymous):

Oh sorry, that should be \[-(x-2)^2+12\]

OpenStudy (anonymous):

The thing is, now that the value of a is negative, the parabola opens downward, so the vertex represents the maximum value of the function.

OpenStudy (vortish):

ok I understand the lanquage of it it is the doing of it i dont understand

OpenStudy (anonymous):

You mean how to come up with the numbers in order to rewrite f(x)?

OpenStudy (vortish):

yes how to work the problem i get confused with steps

OpenStudy (anonymous):

Okay. Is it alright if I try to explain the process using variables (a, b, and c) or would you prefer something more concrete (actual number coefficients)?

OpenStudy (vortish):

a b and c work for me

OpenStudy (anonymous):

Alright \[\begin{align*}y&=ax^2+bx+c\\ y&=a\left(x^2+\frac{b}{a}x+\frac{c}{a}\right)\\ y&=a\left(x^2+\frac{b}{a}x+\left(\frac{b^2}{(2a)^2}-\frac{b^2}{(2a)^2}\right)+\frac{c}{a}\right) \end{align*}\] Do you follow so far?